please answer and show work. Thanks!A pellet of mass 2.51 g from a pellet-gun is fired horozontally at a tin can of mass 185 g poised on a post at a height of 1.
Answer #1

First, find the horizontal velocity of the can after the collision

During collision momentum is conserved

Initial momentum = Final momentum

mpellet * Upellet +mcan * Ucan = m pellet * Upellet + mcan * Ucan

0.00251kg*91m/s+0.185kg*0m/s=0.00251kg*71m/s+0.185kg*v_{can}

0.00251 * 91 + 0.185 *0= 0.00251 * 71 +0.185 * v can

0.22841 = 0.17821 + 0.185 * Ucan

0.22841 -0.17821 = 0.185 * Ucan

\frac{0.22841-0.17821}{0.185}=v_{can}

{\color{Red} v_{can}=0.27135m/s}

-----------------------------------

Consider the vertical motion of the can

Use formula s= ut + -at

h=u_{y}t+\frac{1}{2}gt^{2}

1.2m=0m/s*t+0.5*9.81m/s^{2}*t^{2}

1.2=0*t+0.5*9.81*t^{2}

1.2=0.5*9.81*t^{2}

\sqrt{\frac{1.2}{0.5*9.81}}=t

{\color{Red} t=0.495s}

================

Consider the horizontal motion of the can

There is no acceleration in horizontal direction, so horizontal velocity remains constant.

Horizontal velocity = Horizontal distance/time

v_{can}=\frac{d}{t}

d=v_{can}*t

d=0.27135m/s*0.495s

ANSWER: {\color{Red} d=0.13432m}

=========================